### v2

@@ 4 Results
In Fig. 1, we present a Mollweide projection map displaying the distribution of our final data set of 8504 stars in galactic coordinates (6058 K-Giants and 2446 RR-Lyraes). To enhance the visual representation, the map has been smoothed using Gaussian functions with a full width at half-maximum (FWHM) of 30°. The colour bar represents the density contrast, indicating relative density variation from its mean value across the sample. The past 1 Gyr orbit of the centre of mass of the MCS is shown with a magenta line, computed with the gala package using the MilkyWayPotential (Bovy, 2015). The blue and red dotted lines represent the Pisces (Chandra et al., 2023b) and Virgo overdensities (Perottoni et al., 2022), respectively. The green lines are the polynomials adjusted by Chandra et al. (2023b) that limit the Pisces Plume. The black dots represent the members of the Magellanic stream taken from Chandra et al. (2023a). The dark blue region corresponds to the masked region representing the galactic plane.
Refer to captionRefer to caption
@@
Figure 1: Density distribution of K-Giant and RR-Lyrae variables (mollweide projection map) with $6030<R_{gal}<$ 60 kpc (top panel) and with $10060<R_{gal}<$ 100 kpc (bottom panel). The data is smoothed using a Full-Width at Half Maximum (FWHM) of $30^{\circ}$. The magenta line represents the past orbit of the MCS centre of mass. The blue and red dashed lines represent the Pisces and Virgo overdensities, respectively. The black dots represent the members of the Magellanic Stream. The green lines are the Pisces Plume.
@@
It can be observed, two distinct regions of overdensities. The first one, located in the northern hemisphere, with a longitude range between 225° and 315°, is associated with the collective response, also known as the global response. On the other hand, the southern feature appears to cover a larger area ($-30^{\circ}<l<130^{\circ}$) and exhibits significant prominence at a longitude of $50^{\circ}$ and a latitude of $-26^{\circ}$ approximated. This overdensity is associated with the local wake. A more minor component is also present in the northern hemisphere, within a longitude range of 30° to 90°. It appears separate from the southern component due to the masking of the galactic plane, implemented to prevent contamination. The intensity of the wake is greater than the one of the collective response. The ratio between the counts per pixel of the wake at $l=50^{\circ}$, $b=-26^{\circ}$ and the counts per pixel of the collective response at $l=279^{\circ}$, $b=24^{\circ}$ (coordinates of the highest overdensity of the collective response) is 1.29 considering the complete data set. As one can see, the CM of MCS past orbit is located over the local wake, and the deviations could be signalling the effect of the DM mass of the wake according to the results of Foote et al. (2023).
@@
In the upper panel of Fig. 1, an inner region of the halo is shown (sources between 30 and 60 kpc from the galactic centre). It can be noted that both the wake and the collective do not correspond to either the Pisces or the Virgo overdensities. In previous works (Belokurov et al., 2019; Conroy et al., 2021) it was noticed that there exists a subregion produced by the Magellanic Clouds called the Pisces Plume. The southern overdensity that we identified as the wake does not fall within this region.
@@
On the other hand, in the lower panel, we plot the outer region of the halo, between 60 and 100 kpc, region studied by Belokurov et al. (2019) and Conroy et al. (2021). Once again, the collective does not belong to the Virgo overdensity. Nevertheless, the global response could be truncated due to the masking of the galactic plane and the Sagittarius stream. However, part of the wake lies on the edge of the Pisces overdensity, and the maximum of the wake is indeed located in the Pisces Plume. Additionally, it has been verified that none of the catalogued points belonging to the Magellanic Stream (Chandra et al., 2023a) are found in our dataset.
@@
Comparing our results with the ones obtained by Conroy et al. (2021), we observe some slight differences in the locations of the overdensities. Specifically, our sample’s maximum southern overdensity is slightly displaced further north. Similarly, the maximum northern overdensity in our sample is slightly shifted towards the south and east. In particular, we have also compared only our K-Giant sample results with the final public catalogue developed by Conroy et al. (2021), the coordinates of the maximum overdensities of the local wake and the collective response for our data set are ($l=52^{\circ}$, $b=-24^{\circ}$) and ($l=275^{\circ}$, $b=25^{\circ}$) respectively, meanwhile for Conroy’s data are approximated ($l=49^{\circ}$, $b=-54^{\circ}$) for the wake and ($l=326^{\circ}$, $b=54^{\circ}$) for the collective. Our definition of the local wake is larger than the one proposed by Conroy et al. (2021), and the ratio between the counts per pixel of the wake and the counts per pixel of the collective response at each maximum is 1.5 for our data and 1.33 for Conroy’s data-set. However, if we consider only K-giants located within $10060<R_{gal}<$ 100 kpc, we successfully replicated the wake’s position, with its peak occurring at $l=57^{\circ}$, $b=-51^{\circ}$. Moreover, when we compared our map with the simulations presented in Conroy et al. (2021), we observed quite an agreement regarding the positions of the overdensities.
@@
In Fig. 2 we plot the superficial overdensity of the wake along with the past orbit of the CM (magenta line), in a new coordinate frame, namely orbit frame (see appendix A.2). In this frame, the plane $x^{*}-y^{*}$ contains the CM orbit and the $z^{*}$-axis is perpendicular to the CM orbit. The origin of this new coordinate frame is the current location of the CM, and the $x^{*}$-axis is coincident with the direction of the DM subhalo mass velocity. As one can see, the wake is located in $x^{*}<0$ and it moves towards the perturber. In particular, the lower panel of Fig. 2 is in good agreement with the results presented in Fig. 1 of Buschmann et al. (2018).
Refer to captionRefer to caption
@@
Figure 2: Overdensity as a function of the position. Magenta line: past orbit of the CM, magenta star: current position of the CM, pink arrow: velocity of the CM. Red arrow: mean velocity of the wake/collective, green arrow: mean velocity of a bin of 15 kpc. Cyan star: position of the galactic centre. Top panel: orbit plane with $|z^{*}|<10\,\rm kpc$; bottom panel: $x^{*}-z^{*}$-plane (perpendicular to the orbit plane).
@@
The maximum value of the density is approximated at $x^{*}=-48.97$ kpc, $y^{*}=55.65$ kpc and $z^{*}=-7.94$ kpc in the new orbit frame. Using the stars in a $10$ kpc neighbourhood (177 stars) we computed the velocity dispersion resulting in $(48\pm 3)$ km/s. The characterisation of its complete stellar population and dynamical properties will be addressed in a forthcoming work.
@@
The Gaia satellite data and spectroscopic surveys have unveiled the structure, composition, and formation history of the Milky Way’s stellar halo in recent years (Helmi et al., 2018; Belokurov et al., 2019; Kruijssen et al., 2020; Callingham et al., 2022). Satellite galaxies, globular clusters, stars, and streams are now associated with different halo components, which constitute the remains of our galaxy’s past merger events.
@@
In order to study the possible origin of the stellar populations of the wake and collective response we have performed an $E$ vs $L_{z}$ diagram, see Fig. 3. We have taken the same coordinate system convention as the one adopted by Callingham et al. (2022). In this diagram we have discriminated the different known mergers of the Milky Way (colour points, data extracted from Callingham et al. (2022)), the LMC (star), the SMC (squared) and the wake and collective response (cyan and magenta regions respectively). As one can see, both wake and collective are extended regions in the diagram without a defined sign of $L_{z}$ but limited in energy, however, it is not in the range of the Gaia-Sausage-Enceladus energy. The mean values for the wake are $E=-0.703\times 10^{5}\,\rm{km}^{2}/\rm{s}^{2}$, $L_{z}=890.433\,\rm{km\,kpc\,s}^{-1}$, meanwhile for the collective are $E=-0.716\times 10^{5}\,\rm{km}^{2}/\rm{s}^{2}$, $L_{z}=902.391\,\rm{km\,kpc\,s}^{-1}$.
@@
Figure 3: $E$-$L_{z}$ diagram. The mergers data were taken from Callingham et al. (2022). Cyan region: wake, magenta region: collective response.
We can interpret the findings presented in Fig. 3 as evidence that the Magellanic clouds are dynamically impacting the current stellar population of the Milky Way’s stellar halo. This effect is independent of the stars’ origin and impacts populations that were accreted with Sequoia, Saggitarius, Helmi, ED-3-4-5-6, Typhon, L-RL64 (Dodd et al., 2023) and even residing at the disc (Petersen & Peñarrubia, 2020). Notice that the mergers are of older origin, consequently, the recent impact of the LMC affects dynamically all these stellar populations. This is according to the idea of a first-infall scenario of the LMC’s (Besla et al., 2007; Kallivayalil et al., 2013; Sheng et al., 2024).
@@
Next, we performed the statistical analysis in order to obtain the dark matter subhalo mass of the centre of mass and therefore the DM subhalo mass of the LMC. The radius for the region of interest was fixed at $R=100\,{\rm kpc}$, the subhalo velocity $v_{s}$ was fixed at $314.23$ km/s (van der Marel et al., 2002; Martínez-Delgado et al., 2019) and $n_{0}$ was obtained from the reduced observational data described in the previous section. For the velocity dispersion, we have performed a statistical analysis of the data and obtained the velocity standard deviation in each axis. For the analysis using Eqs. (2), we have considered the larger component of $\bar{v}_{0}$ in the calculations. We have used the described density profile in Section 3 and we present our results in Fig. 4, along with the DM subhalo mass estimation of the LMC published in the literature (Watkins et al., 2024; Koposov et al., 2023; Shipp et al., 2021; Vasiliev et al., 2021; Erkal et al., 2019; Peñarrubia et al., 2016) (using different method indicated with colours) along with our results (last three values, with their correspondent statistical errors).
@@
The dark matter subhalo mass was computed by using the space distribution function (case (a)), the phase-space distribution function of Eq.(3) (case (b)), of Eq.(2) (case (c)). Our results are consistent despite the distribution function used in the analysis. However, the fit obtained using only the space data is slightly higher than the $6D$ analysis results. Furthermore, our findings agree within $3\sigma$ with the literature (Vasiliev, 2023).
@@
Figure 4: Dark matter subhalo mass estimation of the LMC. Orange line: kinematic estimation from Milky Way satellites. Red lines: estimation from stellar streams. Green line: estimation based on momentum balance in the Local Group. Blue lines: our results obtained from the likelihood analysis, case (a): $3D$ data, $M_{LMC}=2.289^{+0.260}_{-0.240}\times 10^{11}\rm{M}_{\odot}$; case (b): $6D$ data (Eq.(3)), $M_{LMC}=1.787^{+0.072}_{-0.069}\times 10^{11}\rm{M}_{\odot}$; case (c): $6D$ data (Eq.(2)), $M_{LMC}=1.686^{+0.071}_{-0.072}\times 10^{11}\rm{M}_{\odot}$.
@@
We have also performed the statistical analysis ($6D$ Eq.(2)) using only the data that have measured radial velocity (Gaia Collaboration et al., 2022; Ahn et al., 2012; Wang et al., 2022). We have found a LMC subhalo mass of $M_{LMC}=1.594^{+0.203}_{-0.196}\times 10^{11}\rm{M}_{\odot}$, that is in good agreement with our results.
@@ 5 Conclusions
In this work, we employed the recently published Gaia Data Release 3 which improves the precision of proper motions along with the Segue catalogue (Ahn et al., 2012) and the one provided by Wang et al. (2022). This enabled us to extend the K-Giants catalogue originally provided by Conroy et al. (2021), and also construct a catalogue for RR-Lyrae stars both in 6D data. We reproduced the previously published results and identified the overdensities associated with a wake and the collective response using these two Halo tracers. A notable finding of this study is the extension of the southern overdensity towards lower galactocentric distances, that is, between 30 and 100 kpc. Moreover, we were able to show that the southern overdensity, identified as the wake, trails the centre of mass of Magellanic Clouds (see Fig. 2).
We have confirmed that the Pisces plume overdensity, described in Belokurov et al. (2019), is associated with the wake of the Magellanic Clouds in the outer regions of the Milky Way’s halo. Furthermore, we have discovered that the overdensity on the halo’s stellar population, caused by the Magellanic Clouds’ wake and the global response, affects the stars in the Milky Way’s halo, regardless of which past merger event they were accreted.
As per the theoretical proposal made by Buschmann et al. (2018), we were able to estimate the mass of the LMC dark matter subhalo for the first time by using Gaia observational data. We found a reliable estimation of the dark matter halo surrounding the LMC by performing two different analyses, using only the space distribution data and using both, the phase and space data. Considering a relationship between the Large and Small Magellanic Clouds’ mass, our study has successfully determined the mass of the dark matter subhalo of the larger cloud. Even more, our findings are in agreement with prior results (Correa Magnus & Vasiliev, 2021; Koposov et al., 2023; Shipp et al., 2021; Vasiliev et al., 2021; Erkal et al., 2019; Peñarrubia et al., 2016), within $3\sigma$. This consistency with previous studies indicates the reliability of our methodology. Additionally, this method gives competitive errors compared to different mass determination methods. It is important to point out that the errors mentioned were calculated using the assumptions of a Plummer profile for a spherical subhalo. However, if a more complex multiparameter model such as an ellipsoidal NFW model is used, it is anticipated that the errors will increase.

### v3

@@ 4 Results
In Fig. 1, we present a Mollweide projection map displaying the distribution of our final dataset of 8504 stars in galactic coordinates (6058 K giants and 2446 RR Lyraes). To enhance the visual representation, the map has been smoothed using Gaussian functions with a full width at half maximum (FWHM) of $30^{\circ}$. The colour bar represents the density contrast, indicating the relative density variation from its mean value across the sample. The past 1 Gyr orbit of the CM of the MCS is shown with a magenta line, computed with the gala package using the MilkyWayPotential (Bovy, 2015). The dotted blue and red lines represent the Pisces (Chandra et al., 2023b) and Virgo overdensities (Perottoni et al., 2022), respectively. The green lines are the polynomials adjusted by Chandra et al. (2023b) that limit the Pisces Plume. The black dots represent the members of the Magellanic Stellar Stream taken from Chandra et al. (2023a). The dark blue region corresponds to the masked region representing the galactic plane.
@@
Figure 1: Density distribution of K-giant and RR-Lyrae variables (Mollweide projection map) with $6030<R_{gal}<$ 60 kpc (top panel) and $10060<R_{gal}<$ 100 kpc (bottom panel). The data was smoothed using an FWHM of $30^{\circ}$. The magenta line represents the past orbit of the MCS CM. The dashed blue and red lines represent the Pisces and Virgo overdensities, respectively. The black dots represent the members of the Magellanic Stellar Stream. The green lines are the Pisces Plume.
@@
Two distinct regions of overdensities can be observed. The first one, located in the northern hemisphere, with a longitude range between 225° and 315°, is associated with the collective response, also known as the global response. On the other hand, the southern feature appears to cover a larger area ($-30^{\circ}<l<130^{\circ}$) and exhibits significant prominence at a longitude of $50^{\circ}$ and a latitude of approximately $-26^{\circ}$. This overdensity is associated with the local wake. A more minor component is also present in the northern hemisphere, within a longitude range of 30° to 90°. It appears separate from the southern component due to the masking of the galactic plane, implemented to prevent contamination. The intensity of the wake is greater than that of the collective response. The ratio between the counts per pixel of the wake at $l=50^{\circ}$, $b=-26^{\circ}$ and the counts per pixel of the collective response at $l=279^{\circ}$, $b=24^{\circ}$ (the coordinates of the highest overdensity of the collective response) is 1.29, considering the complete dataset. As one can see, the CM of the MCS past orbit is located over the local wake, and the deviations could be signalling the effect of the DM mass of the wake according to the results of Foote et al. (2023).
@@
In the upper panel of Fig. 1, an inner region of the halo is shown (sources between 30 and 60 kpc from the galactic centre). It can be noted that both the wake and the collective do not correspond to either the Pisces or the Virgo overdensities. In previous works (Belokurov et al., 2019; Conroy et al., 2021), it was noticed that there exists a subregion produced by the Magellanic Clouds called the Pisces Plume. The southern overdensity that we identified as the wake does not fall within this region.
@@
On the other hand, in the lower panel, we plot the outer region of the halo, between 60 and 100 kpc, a region studied by Belokurov et al. (2019) and Conroy et al. (2021). Once again, the collective does not belong to the Virgo overdensity. Nevertheless, the global response could be truncated due to the masking of the galactic plane and the Sagittarius stream. However, part of the wake lies on the edge of the Pisces overdensity, and the maximum of the wake is indeed located in the Pisces Plume. Additionally, it has been verified that none of the catalogued points belonging to the Magellanic Stellar Stream (Chandra et al., 2023a) are found in our dataset.
@@
Comparing our results with the ones obtained by Conroy et al. (2021), we observe some slight differences in the locations of the overdensities. Specifically, our sample’s maximum southern overdensity is slightly displaced further north. Similarly, the maximum northern overdensity in our sample is slightly shifted towards the south and east. In particular, we have also compared only our K-giant sample results with the final public catalogue developed by Conroy et al. (2021). The coordinates of the maximum overdensities of the local wake and the collective response for our dataset are ($l=52^{\circ}$, $b=-24^{\circ}$) and ($l=275^{\circ}$, $b=25^{\circ}$), respectively; meanwhile, for Conroy’s data they are approximately ($l=49^{\circ}$, $b=-54^{\circ}$) for the wake and ($l=326^{\circ}$, $b=54^{\circ}$) for the collective. Our definition of the local wake is larger than the one proposed by Conroy et al. (2021), and the ratio between the counts per pixel of the wake and the counts per pixel of the collective response at each maximum is 1.5 for our data and 1.33 for Conroy’s data-set. However, if we consider only K giants located within $10060<R_{gal}<$ 100 kpc, we successfully replicated the wake’s position, with its peak occurring at $l=57^{\circ}$, $b=-51^{\circ}$. Moreover, when we compare our map with the simulations presented in Conroy et al. (2021), we observe quite an agreement regarding the positions of the overdensities.
@@
In Fig. 2, we plot the superficial overdensity of the wake along with the past orbit of the CM (magenta line), in a new coordinate frame; namely, the orbit frame (see appendix A.2). In this frame, the plane $x^{*}-y^{*}$ contains the CM orbit and the $z^{*}$ axis is perpendicular to the CM orbit. The origin of this new coordinate frame is the current location of the CM, and the $x^{*}$ axis is coincident with the direction of the DM subhalo mass velocity. As one can see, the wake is located in $x^{*}<0$ and it moves towards the perturber. In particular, the lower panel of Fig. 2 is in good agreement with the results presented in Fig. 1 of Buschmann et al. (2018).
@@
Figure 2: Overdensity as a function of the position. Magenta line: past orbit of the CM. Magenta star: current position of the CM. Pink arrow: velocity of the CM. Red arrow: mean velocity of the wake or collective. Green arrow: mean velocity of a bin of 15 kpc. Cyan star: position of the galactic centre. Top panel: Orbit plane with $|z^{*}|<10\,\rm kpc$. Bottom panel: $x^{*}-z^{*}$ plane (perpendicular to the orbit plane).
@@
The maximum value of the density is approximately $x^{*}=-48.97$ kpc, $y^{*}=55.65$ kpc, and $z^{*}=-7.94$ kpc in the new orbit frame. Using the stars in a $10$ kpc neighbourhood (177 stars), we computed the velocity dispersion, resulting in $(48\pm 3)$ km/s. The characterisation of its complete stellar population and dynamical properties will be addressed in a forthcoming work.
@@
The Gaia satellite data and spectroscopic surveys have unveiled the structure, composition, and formation history of the MW’s stellar halo in recent years (Helmi et al., 2018; Belokurov et al., 2019; Kruijssen et al., 2020; Callingham et al., 2022). Satellite galaxies, globular clusters, stars, and streams are now associated with different halo components, which constitute the remains of our galaxy’s past merger events.
@@
In order to study the possible origin of the stellar populations of the wake and collective response, we have made an $E$ vs $L_{z}$ diagram (see Fig. 3). We have taken the same coordinate system convention as the one adopted by Callingham et al. (2022). In this diagram, we have discriminated the different known mergers of the MW (colour points, data extracted from Callingham et al. (2022)), the LMC (star), the SMC (squared), and the wake and collective response (cyan and magenta regions, respectively). As one can see, both the wake and the collective are extended regions in the diagram without a defined sign of $L_{z}$ but limited in energy; however, it is not in the range of the Gaia-Sausage-Enceladus energy. The mean values for the wake are $E=-0.703\times 10^{5}\,\rm{km}^{2}/\rm{s}^{2}$ and $L_{z}=890.433\,\rm{km\,kpc\,s}^{-1}$; meanwhile, for the collective, they are $E=-0.716\times 10^{5}\,\rm{km}^{2}/\rm{s}^{2}$ and $L_{z}=902.391\,\rm{km\,kpc\,s}^{-1}$.
@@
Figure 3: $E$-$L_{z}$ diagram. The mergers data were taken from Callingham et al. (2022). Cyan region: wake. Magenta region: collective response.
It should be noted that the stellar population of the MW’s Halo was accreted by mergers of older origin such as Sequoia, Saggitarius, Helmi, ED-3-4-5-6, Typhon, and L-RL64 (Dodd et al., 2023). Consequently, the recent impact of the LMC could dynamically affect all these stellar populations; however, more studies are needed.
@@
Next, we performed a statistical analysis in order to obtain the DM subhalo mass of the CM, and therefore the DM subhalo mass of the LMC. The radius for the region of interest was fixed at $R=100\,{\rm kpc}$, the subhalo velocity, $v_{s}$, was fixed at $314.23$ km/s (van der Marel et al., 2002; Martínez-Delgado et al., 2019), and $n_{0}$ was obtained from the reduced observational data described in the previous section. For the velocity dispersion, we performed a statistical analysis of the data and obtained the velocity standard deviation in each axis. For the analysis using Eqs. (2), we considered the larger component of $\bar{v}_{0}$ in the calculations. We used the described density profile in Section 3 and we present our results in Fig. 4, along with the DM subhalo mass estimation of the LMC published in the literature (Watkins et al., 2024; Koposov et al., 2023; Shipp et al., 2021; Vasiliev et al., 2021; Erkal et al., 2019; Peñarrubia et al., 2016) (using a different method, indicated with colours) along with our results (the last three values, with their corresponding statistical errors).
@@
The DM subhalo mass was computed by using the space distribution function (case (a)) and the phase-space distribution function of Eq.(3) (case (b)) and Eq.(2) (case (c)). Our results are consistent, despite the distribution function used in the analysis. However, the fit obtained using only the space data is slightly higher than the $6D$ analysis results. Furthermore, our findings agree within $3\sigma$ with the literature (Vasiliev, 2023).
@@
Figure 4: Dark matter subhalo mass estimation of the LMC. Orange line: kinematic estimation from MW satellites. Red lines: estimation from stellar streams. Green line: estimation based on momentum balance in the Local Group. Blue lines: our results obtained from the likelihood analysis. Case (a): $3D$ data, $M_{LMC}=2.289^{+0.260}_{-0.240}\times 10^{11}\rm{M}_{\odot}$. Case (b): $6D$ data (Eq.(3)), $M_{LMC}=1.787^{+0.072}_{-0.069}\times 10^{11}\rm{M}_{\odot}$. Case (c): $6D$ data (Eq.(2)), $M_{LMC}=1.686^{+0.071}_{-0.072}\times 10^{11}\rm{M}_{\odot}$. The error bars of our results are purely statistical and based on a restricted one-parameter model.
@@
We also performed the statistical analysis ($6D$ Eq.(2)) using only the data with measured radial velocity (Gaia Collaboration et al., 2023; Ahn et al., 2012; Wang et al., 2022). We found an LMC subhalo mass of $M_{LMC}=1.594^{+0.203}_{-0.196}\times 10^{11}\rm{M}_{\odot}$, which is in good agreement with our results.
@@ 5 Conclusions
In this work, we employed the recently published Gaia Data Release 3, which improves the precision of proper motions along with the Segue catalogue (Ahn et al., 2012) and the one provided by Wang et al. (2022). This enabled us to extend the K-giant catalogue originally provided by Conroy et al. (2021), and also to construct a catalogue for RR-Lyrae stars, both in 6D data. We reproduced the previously published results and identified the overdensities associated with a wake and the collective response using these two halo tracers. A notable finding of this study is the extension of the southern overdensity towards lower galactocentric distances; that is, between 30 and 100 kpc. Moreover, we were able to show that the southern overdensity, identified as the wake, trails the CM of Magellanic Clouds (see Fig. 2).
We have confirmed that the Pisces plume overdensity, described in Belokurov et al. (2019), is associated with the wake of the Magellanic Clouds in the outer regions of the MW’s halo. Furthermore, we have discovered that the overdensity on the halo’s stellar population, caused by the Magellanic Clouds’ wake and the global response, affects the stars in the MW’s halo, regardless of which past merger event they were accreted from.
As per the theoretical proposal made by Buschmann et al. (2018), we were able to estimate the mass of the LMC DM subhalo for the first time by using Gaia observational data. We found a reliable estimation of the DM halo surrounding the LMC by performing two different analyses, using only the space distribution data and using both the phase and space data. Considering a relationship between the Large and Small Magellanic Clouds’ masses, our study has successfully determined the mass of the DM subhalo of the larger cloud. Even more, our findings are in agreement with prior results (Correa Magnus & Vasiliev, 2021; Koposov et al., 2023; Shipp et al., 2021; Vasiliev et al., 2021; Erkal et al., 2019; Peñarrubia et al., 2016), within $3\sigma$. This consistency with previous studies indicates the reliability of our methodology. Additionally, this method gives competitive errors compared to different mass determination methods. It is important to point out that the errors mentioned were calculated using the assumptions of a Plummer profile for a spherical subhalo. However, if a more complex multi-parameter model such as an ellipsoidal Navarrro, Frenk &\&& White (NFW) model were used, it is anticipated that the errors would increase.